# Neural Operators Learn the Local Physics of Magnetohydrodynamics

### v1

@@ 3.2 Enhanced Loss Function for Ideal MHD
$\mathcal{L}_{TVD}(\{\mathbb{U}_{i}\}_{i=1}^{B}):=\sum_{i=1}^{B}\sum_{n=1}^{N_{t}-1}\lfloor TV(\mathbb{\tilde{U}}_{i,n+1})-TV(\mathbb{U}_{i,n})\rfloor_{+}^{2}$ (12)
@@
$\Rightarrow 0=\frac{\partial\nabla\cdot\mathbb{B}}{\partial t}=\nabla\cdot\Big{(}-\frac{\partial\mathbb{G}}{\partial x}-\frac{\partial\mathbb{F}}{\partial y}\Big{)}_{\mathbb{B}}$ (13)
@@
$\mathcal{L}_{div}(\{\mathbb{U_{i}}\}_{i=1}^{B}):=\sum_{i=1}^{B}\sum_{j=1}^{N_{t}-1}\Bigg{\lfloor}\frac{\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)_{\mathbb{B}}\|_{2}^{2}-\theta_{div}}{|\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)_{\mathbb{B}}\|_{2}^{2}-\theta_{div}|}\Bigg{\rfloor}_{+}\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)_{\mathbb{B}}\|_{2}^{2}$ (14)
@@
($\mathcal{L}_{\infty}(\{\mathbb{U_{i}}\}_{i=1}^{B}):=\sum_{i=1}^{B}\sum_{n=1}^{N_{t}-1}\sum_{l=1}^{N_{u}}\sup_{j,k}\Big{(}(\mathbb{U}_{i,n+1,j,k,l}-\mathbb{U}_{i,n,j,k,l})-\Delta\mathbb{R}(\mathbb{U}_{i,n,j,k,l};\theta)\Big{)}$ (15)
@@ 3.3 Overall Algorithm
Input: Dataset $\mathcal{U}=((U_{b,i,j,k,m}),(\Delta t_{b,i}))$
@@
         $\tilde{U}_{-\tilde{j}}\leftarrow$ roll $U_{b,i,\cdot,\cdot,m}$ by $\tilde{j}$ in the third index for $\tilde{j}=-q,\dots,p+1$
         $\tilde{U}^{-\tilde{k}}\leftarrow$ roll $U_{b,i,\cdot,\cdot,m}$ by $\tilde{k}$ in the fourth index for $\tilde{k}=-q,\dots,p+1$
         $U^{l}\leftarrow$ concatenate along fifth index $(\tilde{U}_{-p},\dots,\tilde{U}_{q})$
         $U^{r}\leftarrow$ concatenate along fifth index $(\tilde{U}_{-p-1},\dots,\tilde{U}_{q-1})$
         $U^{t}\leftarrow$ concatenate along fifth index $(\tilde{U}^{-p},\dots,\tilde{U}^{q})$
         $U^{b}\leftarrow$ concatenate along fifth index $(\tilde{U}^{-p-1},\dots,\tilde{U}^{q-1})$triangleright Thus, the concatenated function is now a 8(p+q)-dimensional vector-valued function
         $G^{l}\leftarrow$ concatenate along fifth index $(G_{\rho}(U^{l};\theta_{\rho}^{G}),G_{\mathbb{u}}(U^{l};\theta_{\mathbb{u}}^{G}),G_{\mathbb{B}}(U^{l};\theta_{\mathbb{B}}^{G}),G_{E}(U^{l};\theta_{E}^{G}))$
         $G^{r}\leftarrow$ concatenate along fifth index $(G_{\rho}(U^{r};\theta_{\rho}^{G}),G_{\mathbb{u}}(U^{r};\theta_{\mathbb{u}}^{G}),G_{\mathbb{B}}(U^{r};\theta_{\mathbb{B}}^{G}),G_{E}(U^{r};\theta_{E}^{G}))$
         $F^{t}\leftarrow$ concatenate along fifth index $(F_{\rho}(U^{t};\theta_{\rho}^{F}),F_{\mathbb{u}}(U^{t};\theta_{\mathbb{u}}^{F}),F_{\mathbb{B}}(U^{t};\theta_{\mathbb{B}}^{F}),F_{E}(U^{t};\theta_{E}^{F}))$
         $F^{b}\leftarrow$ concatenate along fifth index $(F_{\rho}(U^{b};\theta_{\rho}^{F}),F_{\mathbb{u}}(U^{b};\theta_{\mathbb{u}}^{F}),F_{\mathbb{B}}(U^{b};\theta_{\mathbb{B}}^{F}),F_{E}(U^{b};\theta_{E}^{F}))$
         $\Delta\mathbb{R}(\mathbb{U}_{b,i,\cdot,\cdot,\cdot};\theta)\leftarrow$ $\frac{\Delta t_{b,i}}{\Delta x}\Big{[}G^{l}-G^{r}\Big{]}+\frac{\Delta t_{b,i}}{\Delta y}\Big{[}F^{t}-F^{b}\Big{]}$
         $\mathcal{L}_{tm}(Batch)\leftarrow\sum_{b=1}^{B}\sum_{i=1}^{N_{t}-1}\|U_{b,i+1,\cdot,\cdot,\cdot}-U_{b,i,\cdot,\cdot,\cdot}-\Delta\mathbb{R}(\mathbb{U}_{b,i,\cdot,\cdot,\cdot};\theta)\|_{2}^{2}$triangleright B is batch size
@@
         $\mathcal{L}_{consi}(Batch)\leftarrow\sum_{b=1}^{B}\sum_{i=1}^{N_{t}}\Big{(}\|\tilde{G}(V^{p+q}_{b,i,\cdot,\cdot,\cdot};\theta)-G(V^{p+q}_{b,i,\cdot,\cdot,\cdot})\|_{2}^{2}+\|\tilde{F}(V^{p+q}_{b,i,\cdot,\cdot,\cdot};\theta)-F(V^{p+q}_{b,i,\cdot,\cdot,\cdot})\|_{2}^{2}\Big{)}$triangleright Each of $\tilde{G}$ and $\tilde{F}$ represents concatenated numerical fluxes, and $G$ and $F$ are the physical fluxes.
         $\mathcal{L}_{TVD}(Batch)\leftarrow\sum_{b=1}^{B}\sum_{i=1}^{N_{t}-1}\|\lfloor TV(U_{b,i,\cdot,\cdot,\cdot}+\Delta\mathbb{R}(\mathbb{U}_{b,i,\cdot,\cdot,\cdot};\theta))-TV(U_{b,i,\cdot,\cdot,\cdot})\rfloor_{+}\|_{2}^{2}$
         $\mathcal{L}_{\infty}(Batch)\leftarrow\sum_{b=1}^{B}\sum_{i=1}^{N_{t}-1}\sum_{l=1}^{N_{u}}\sup_{j,k}\|U_{b,i+1,j,k,l}-U_{b,i,j,k,l}-\Delta\mathbb{R}(\mathbb{U}_{b,i,j,k,l};\theta)\|_{2}^{2}$
         $\mathcal{L}_{div}(Batch)\leftarrow\sum_{i=1}^{B}\sum_{j=1}^{N_{t}-1}\Bigg{\lfloor}\frac{\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)_{\mathbb{B}}\|_{2}^{2}-\theta_{div}}{|\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)_{\mathbb{B}}\|_{2}^{2}-\theta_{div}|}\Bigg{\rfloor}_{+}\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)_{\mathbb{B}}\|_{2}^{2}$
@@ 4.1 Training Dataset and Architecture of Flux Neural Operator
One-dimensional case For the one-dimensional problem, we constructed the initial conditions using Gaussian random fields. Specifically, we set the initial values for density and the $B_{x}$ to 1, and specified the velocity and magnetic field components as vectors discretized to size 256, derived from Gaussian random fields with a covariance of $k(x,y)=e^{-100(x-y)^{2}}$. To ensure the well-posedness of the problem, each component of the velocity and magnetic field was scaled by 0.3. Given that the $\gamma$ value was set to 2.0, the initial value of the energy is determined by the following equation: $E=1.0+\frac{1}{2}\rho\mathbb{u}^{2}+\frac{1}{2}\|\mathbb{B}\|^{2}$. Based on these initial conditions, we solved the problem using the WENO-JS scheme for spatial order 5 and the RK method of order 4 for temporal discretization. The CFL number was 0.3. Based on this scheme, we divided the solution from $t=0.0$ to $t=0.6$ into 600 segments and then extracted snapshots only from $t=0.1$ to $t=0.6$ to form a dataset consisting of 500 snapshots for each function. We generated 500 tensors representing functions in $C^{\infty}([0.1,0.6]\times[0,1])$ to create the training dataset. In the model architecture, the Neural Operator component fundamentally employs the Fourier neural operator; however, as FNO performs poorly with non-periodic functions, all datasets were composed as periodic functions. We anticipate finding suitable models for non-periodic functions as well, since the Neural Operator component can be replaced with other models. For the one-dimensional problem discussed in this paper, the model’s architecture uses 5 modes, width 64, and depth 3, with output dimensions of 1, 3, 2, and 1 for density, velocity, magnetic field, and energy, respectively. The kernel size of the CNN layer in the Fourier layer is uniformly set to 1.
@@
Two-dimensional case For the two-dimensional problem, initial conditions were generated using two-dimensional Gaussian random fields. We independently sampled the components of velocity and magnetic field vectors from Gaussian random fields. And set the initial density value to $\gamma^{2}$. The Gaussian random fields we used have a power spectrum $P(k)\propto k^{-2.5}$. Unlike the one-dimensional case, in two dimensions, the divergence of the magnetic field must be zero for physically meaningful results and for the numerical scheme to be stable; therefore, the randomly sampled magnetic fields were relaxed using the Poisson equation (Jiang and Wu (1999)). The grid size for the initial conditions was set to 64 along both the x and y axes, and the $\gamma$ value was set to 5/3. The initial energy value will be determined by the following equation: $E=\frac{5}{2}+\frac{1}{2}\rho\mathbb{u}^{2}+\frac{1}{2}\|\mathbb{B}\|^{2}$. Based on these initial conditions, we used the WENO-Z scheme for spatial order 5, temporal discretization of fourth order RK method, and set the CFL numbers to 0.4. Under this scheme, we divided the solution from $t=0.0$ to $t=0.75$ into 150 segments, and then extracted snapshots only from $t=0.25$ to $t=0.75$ to form a dataset consisting of 100 snapshots per function. In this way, we created 100 tensors, each representing a function in $C^{\infty}([0.25,0.75]\times[0,2\pi]^{2})$ for training. The model architecture adopted a 2D FNO for the neural operator component, with architecture hyperparameters including 4 modes each for x and y axes, width 72, and depth 2. The output dimensions for models concerning density, velocity, magnetic field, and energy are 1, 3, 3, and 1, respectively. The kernel size of the CNN layer in the Fourier layer is uniformly set to 1. Both one-dimensional and two-dimensional problems use the GELU activation function. Each neural operator takes concatenated tensors composed of eight tensors, which are shifted circularly from -3 to 4 (where an $i$ shift means $j\rightarrow j+i$) for $G^{l}$ and $F^{t}$, and from -4 to 3 for $G^{r}$ and $F^{b}$ for corresponding indices.
@@ 4.2 Generalization Ability
Short term inference
We conducted experiments on test samples that were sampled from the same distribution as the training dataset and had the same short-term span lengths as the training samples ($t=0.1$ to $0.6$ for the one-dimensional case, $t=0.25$ to $0.75$ for the two-dimensional case). The results for the one-dimensional case are presented in Figure 4 and Table 1, while the results for the two-dimensional case are shown in Figure 5, 6, 14, 7 and Table 2. Tables 1 and 2 provide statistics on the relative $l^{2}$ norm and $l^{\infty}$ norm across 10 test samples. Due to the smaller domain size in the one-dimensional case compared to the two-dimensional case, the changes in the solutions are observed to be more dynamic. In addition, in the one-dimensional case, the solution generated by Flux NO tends to lose stability more easily because the grid size of 256 (compared to 64 in each axis for the two-dimensional case) makes it easier to violate the CFL condition. For the one-dimensional problem, the instability of Flux NO is such that only results up to $t=0.2$ are shown in Table 1. From Table 1, it can be seen that the norms related to velocity are the largest and increase the fastest, which can be attributed to the poorer approximation of the $u_{y}$ and $u_{z}$ components, while $u_{x}$ is well approximated compared to these, as indicated by Figure 4. Consequently, although the training dataset for the one-dimensional problem is richer, consisting of 500 functions compared to 100 for the two-dimensional problem, it is expected to have lower stability and generalization power due to the reasons mentioned above. For the two-dimensional problem, the performance on test samples within the short-term training distribution appears to be satisfactory, as can be seen from Figure 5, 6, 14, 7 and Table 2, indicating that the solution for each component has been well resolved.
@@
Long term inference We conducted experiments over time spans significantly longer than the time domains of the functions used for training, with the initial conditions configured from the same distribution as training dataset. Due to instability issues, long-term experiments were not conducted for the one-dimensional case; instead, they were carried out only for the two-dimensional case, with the results presented in Figure 8, 18, 15, 19 and Table 3. The training functions’ time domain ranged from $t=0.25$ to $t=0.75$, with a fixed time interval of delta $\Delta t=0.005$, allowing for the terminal point to be reached after 100 iterations. However, in our results, using the same time intervals, the iterations extended to much longer times: $t=1.25$ (200 iterations), $t=1.5$ (250 iterations), $t=1.75$ (300 iterations), and $t=2.0$ (350 iterations). As shown in Figures 14 and 15, the error increases almost linearly with the number of iterations, affecting both the global and local characteristics of the solution and leading to progressively increasing discrepancies with the reference, as qualitatively evident in Figures 8, 18 and 19. A notable aspect of the long-term inference for the two-dimensional problem is that, unlike the one-dimensional case, even with many more iterations, the solutions output by Flux NO did not blow up.
@@
High resolution inference We conducted tests on samples at a higher resolution (96 by 96) than the resolution of the training data (64 by 64). Initial conditions were sampled from a Gaussian random field with the same power spectrum as the training data, and references were computed using the WENO-Z scheme with each snapshot spaced by a $\Delta t=0.005$. The test results are presented in Figures 9, 20, and 16; trends in errors, which showed no significant difference from the original resolution, have been omitted. As seen in Figures 9 and 20, while Flux NO captures the global characteristics of the solution, local instabilities gradually appear in the snapshots at $t=1.5$ and $t=2.0$. This instability, similar to that observed in the one-dimensional case, appears to arise from the higher resolution and is somewhat related to the CFL condition. Attempts to more easily satisfy the CFL condition by reducing the interval to $\Delta t=0.0025$ for denser inference were made, yet instabilities still formed. From these experimental results, we hypothesize that as the resolution of vectorized functions increases, the domain dimension of their data distribution expands, making it challenging to approximate the data distribution itself, leading to these outcomes.

### v2

@@ 3.2 Enhanced Loss Function for Ideal MHD
$\mathcal{L}_{TVD}(\{\mathbb{U}_{i}\}_{i=1}^{B}):=\sum_{i=1}^{B}\sum_{n=1}^{N_{t}-1}\lfloor TV(\mathbb{\tilde{U}}_{i,n+1})-TV(\mathbb{U}_{i,n})\rfloor_{+}^{2}.$ (12)
@@
$\Rightarrow 0=\frac{\partial\nabla\cdot\mathbb{B}}{\partial t}=\nabla\cdot\Big{(}-\frac{\partial\mathbb{G}}{\partial x}-\frac{\partial\mathbb{F}}{\partial y}\Big{)}_{\mathbb{B}}.$ (13)
@@
$\mathcal{L}_{div}(\{\mathbb{U_{i}}\}_{i=1}^{B}):=\sum_{i=1}^{B}\sum_{j=1}^{N_{t}-1}\Bigg{\lfloor}\frac{\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)_{\mathbb{B}}\|_{2}^{2}-\theta_{div}}{|\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)_{\mathbb{B}}\|_{2}^{2}-\theta_{div}|}\Bigg{\rfloor}_{+}\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)_{\mathbb{B}}\|_{2}^{2}.$ (14)
@@
($\mathcal{L}_{\infty}(\{\mathbb{U_{i}}\}_{i=1}^{B}):=\sum_{i=1}^{B}\sum_{n=1}^{N_{t}-1}\sum_{l=1}^{N_{u}}\sup_{j,k}\Big{(}(\mathbb{U}_{i,n+1,j,k,l}-\mathbb{U}_{i,n,j,k,l})-\Delta\mathbb{R}(\mathbb{U}_{i,n,j,k,l};\theta)\Big{)}.$ (15)
@@ 3.3 Overall Algorithm
Input: Dataset $\mathcal{U}=((U_{\tilde{b},i,j,k,m}),(\Delta t_{b,i}))$
@@
         $\tilde{U}_{-\tilde{j}}\leftarrow$ roll $U_{\tilde{b},i,\cdot,\cdot,m}$ by $\tilde{j}$ in the third index for $\tilde{j}=-q,\dots,p+1$
         $\tilde{U}^{-\tilde{k}}\leftarrow$ roll $U_{\tilde{b},i,\cdot,\cdot,m}$ by $\tilde{k}$ in the fourth index for $\tilde{k}=-q,\dots,p+1$
         $U^{l}\leftarrow$ concatenate $(\tilde{U}_{-p},\dots,\tilde{U}_{q})$ along fifth index
         $U^{r}\leftarrow$ concatenate $(\tilde{U}_{-p-1},\dots,\tilde{U}_{q-1})$ along fifth index
         $U^{t}\leftarrow$ concatenate $(\tilde{U}^{-p},\dots,\tilde{U}^{q})$ along fifth index
         $U^{b}\leftarrow$ concatenate $(\tilde{U}^{-p-1},\dots,\tilde{U}^{q-1})$ along fifth index $\triangleright$ Thus, the concatenated function is now a 8(p+q)-dimensional vector-valued function
         $G^{l}\leftarrow$ concatenate $(G_{\rho}(U^{l};\theta_{\rho}^{G}),G_{\mathbb{u}}(U^{l};\theta_{\mathbb{u}}^{G}),G_{\mathbb{B}}(U^{l};\theta_{\mathbb{B}}^{G}),G_{E}(U^{l};\theta_{E}^{G}))$ along fifth index
         $G^{r}\leftarrow$ concatenate $(G_{\rho}(U^{r};\theta_{\rho}^{G}),G_{\mathbb{u}}(U^{r};\theta_{\mathbb{u}}^{G}),G_{\mathbb{B}}(U^{r};\theta_{\mathbb{B}}^{G}),G_{E}(U^{r};\theta_{E}^{G}))$ along fifth index
         $F^{t}\leftarrow$ concatenate $(F_{\rho}(U^{t};\theta_{\rho}^{F}),F_{\mathbb{u}}(U^{t};\theta_{\mathbb{u}}^{F}),F_{\mathbb{B}}(U^{t};\theta_{\mathbb{B}}^{F}),F_{E}(U^{t};\theta_{E}^{F}))$ along fifth index
         $F^{b}\leftarrow$ concatenate $(F_{\rho}(U^{b};\theta_{\rho}^{F}),F_{\mathbb{u}}(U^{b};\theta_{\mathbb{u}}^{F}),F_{\mathbb{B}}(U^{b};\theta_{\mathbb{B}}^{F}),F_{E}(U^{b};\theta_{E}^{F}))$ along fifth index
         $\Delta\mathbb{R}(\mathbb{U}_{\tilde{b},i,\cdot,\cdot,\cdot};\theta)\leftarrow$ $\frac{\Delta t_{\tilde{b},i}}{\Delta x}\Big{[}G^{l}-G^{r}\Big{]}+\frac{\Delta t_{\tilde{b},i}}{\Delta y}\Big{[}F^{t}-F^{b}\Big{]}$
         $\mathcal{L}_{tm}(Batch)\leftarrow\sum_{\tilde{b}=1}^{B}\sum_{i=1}^{N_{t}-1}\|U_{\tilde{b},i+1,\cdot,\cdot,\cdot}-U_{\tilde{b},i,\cdot,\cdot,\cdot}-\Delta\mathbb{R}(\mathbb{U}_{\tilde{b},i,\cdot,\cdot,\cdot};\theta)\|_{2}^{2}$triangleright B is batch size
@@
         $\mathcal{L}_{consi}(Batch)\leftarrow\sum_{\tilde{b}=1}^{B}\sum_{i=1}^{N_{t}}\Big{(}\|\tilde{G}(V^{p+q}_{\tilde{b},i,\cdot,\cdot,\cdot};\theta)-G(V^{p+q}_{\tilde{b},i,\cdot,\cdot,\cdot})\|_{2}^{2}+\|\tilde{F}(V^{p+q}_{\tilde{b},i,\cdot,\cdot,\cdot};\theta)-F(V^{p+q}_{\tilde{b},i,\cdot,\cdot,\cdot})\|_{2}^{2}\Big{)}$triangleright Each of $\tilde{G}$ and $\tilde{F}$ represents concatenated numerical fluxes, and $G$ and $F$ are the physical fluxes.
         $\mathcal{L}_{TVD}(Batch)\leftarrow\sum_{\tilde{b}=1}^{B}\sum_{i=1}^{N_{t}-1}\|\lfloor TV(U_{\tilde{b},i,\cdot,\cdot,\cdot}+\Delta\mathbb{R}(\mathbb{U}_{\tilde{b},i,\cdot,\cdot,\cdot};\theta))-TV(U_{\tilde{b},i,\cdot,\cdot,\cdot})\rfloor_{+}\|_{2}^{2}$
         $\mathcal{L}_{\infty}(Batch)\leftarrow\sum_{\tilde{b}=1}^{B}\sum_{i=1}^{N_{t}-1}\sum_{l=1}^{N_{u}}\sup_{j,k}\|U_{\tilde{b},i+1,j,k,l}-U_{\tilde{b},i,j,k,l}-\Delta\mathbb{R}(\mathbb{U}_{\tilde{b},i,j,k,l};\theta)\|_{2}^{2}$
         $\mathcal{L}_{div}(Batch)\leftarrow\sum_{\tilde{b}=1}^{B}\sum_{i=1}^{N_{t}-1}\Bigg{\lfloor}\frac{\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{\tilde{b},i};\theta)_{\mathbb{B}}\|_{2}^{2}-\theta_{div}}{|\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{\tilde{b},i};\theta)_{\mathbb{B}}\|_{2}^{2}-\theta_{div}|}\Bigg{\rfloor}_{+}\|\nabla\cdot\Delta\mathbb{R}(\mathbb{U}_{\tilde{b},i};\theta)_{\mathbb{B}}\|_{2}^{2}$
@@ 4.1 Training Dataset and Architecture of Flux Neural Operator
One-dimensional case For the one-dimensional problem, we constructed the initial conditions using Gaussian random fields. Specifically, we set the initial values for density and the $B_{x}$ to 1, and specified the velocity and magnetic field components as vectors discretized to size 256, derived from Gaussian random fields with a covariance of $k(x,y)=e^{-100(x-y)^{2}}$. To ensure the well-posedness of the problem, each component of the velocity and magnetic field was scaled by 0.3. Given that the $\gamma$ value was set to 2.0, the initial value of the energy is determined by the following equation: $E=1.0+\frac{1}{2}\rho\mathbb{u}^{2}+\frac{1}{2}\|\mathbb{B}\|^{2}$. Based on these initial conditions, we solved the problem using the WENO-JS scheme for spatial order 5 and the RK method of order 4 for temporal discretization. The CFL number was 0.3. Based on this scheme, we divided the solution from $t=0.0$ to $t=0.6$ into 600 segments and then extracted snapshots only from $t=0.1$ to $t=0.6$ to form a dataset consisting of 500 snapshots for each function. We generated 500 tensors representing functions in $C^{\infty}([0.1,0.6]\times[0,1])$ to create the training dataset. In the model architecture, the Neural Operator component fundamentally employs the FNO; however, as FNO performs poorly with non-periodic functions, all datasets were composed as periodic functions. We anticipate finding suitable models for non-periodic functions as well, since the Neural Operator component can be replaced with other models. For the one-dimensional problem discussed in this paper, the model’s architecture uses 5 modes, width 64, and depth 3, with output dimensions of 1, 3, 2, and 1 for density, velocity, magnetic field, and energy, respectively. The kernel size of the CNN layer in the Fourier layer is uniformly set to 1.
@@
Two-dimensional case For the two-dimensional problem, initial conditions were generated using two-dimensional Gaussian random fields. We independently sampled the components of velocity and magnetic field vectors from Gaussian random fields. And set the initial density value to $\gamma^{2}$. The Gaussian random fields we used have a power spectrum $P(k)\propto k^{-2.5}$. Unlike the one-dimensional case, in two dimensions, the divergence of the magnetic field must be zero for physically meaningful results and for the numerical scheme to be stable; therefore, the randomly sampled magnetic fields were relaxed using the Poisson equation (Jiang and Wu (1999)). The grid size for the initial conditions was set to 64 along both the x and y axes, and the $\gamma$ value was set to 5/3. The initial energy value will be determined by the following equation: $E=\frac{5}{2}+\frac{1}{2}\rho\mathbb{u}^{2}+\frac{1}{2}\|\mathbb{B}\|^{2}$. Based on these initial conditions, we used the WENO-Z scheme for spatial order 5, temporal discretization of fourth order RK method, and set the CFL numbers to 0.4. Under this scheme, we divided the solution from $t=0.0$ to $t=0.75$ into 150 segments, and then extracted snapshots only from $t=0.25$ to $t=0.75$ to form a dataset consisting of 100 snapshots per function. In this way, we created 100 tensors, each representing a function in $C^{\infty}([0.25,0.75]\times[0,2\pi]^{2})$ for training. The model architecture adopted a 2D FNO for the neural operator component, with architecture hyperparameters including 4 modes each for the x and y axes, a width of 72, and a depth of 2. These hyperparameters were empirically optimized within the resource constraints available to us. Increasing the maximal frequency did not result in significant performance improvements, which is presumed to be because the physics of ideal MHD inherently propagates information locally. The output dimensions for models concerning density, velocity, magnetic field, and energy are 1, 3, 3, and 1, respectively. The kernel size of the CNN layer in the Fourier layer is uniformly set to 1. Both one-dimensional and two-dimensional problems use the GELU activation function. Each neural operator takes concatenated tensors composed of eight tensors, which are shifted circularly from -3 to 4 (where an $i$ shift means $j\rightarrow j+i$) for $G^{l}$ and $F^{t}$, and from -4 to 3 for $G^{r}$ and $F^{b}$ for corresponding indices.
@@ 4.2 Generalization Ability
Short term inference We conducted experiments on test samples that were sampled from the same distribution as the training dataset and had the same short-term span lengths as the training samples ($t=0.1$ to $0.6$ for the one-dimensional case, $t=0.25$ to $0.75$ for the two-dimensional case). The results for the one-dimensional case are presented in Figure 4 and Table 1, while the results for the two-dimensional case are shown in Figure 5, 6, 14, 7 and Table 2. Tables 1 and 2 provide statistics on the relative $l^{2}$ norm and $l^{\infty}$ norm across 10 test samples. Due to the smaller domain size in the one-dimensional case compared to the two-dimensional case, the changes in the solutions are observed to be more dynamic. In addition, in the one-dimensional case, the solution generated by Flux NO tends to lose stability more easily because the grid size of 256 (compared to 64 in each axis for the two-dimensional case) makes it easier to violate the CFL condition. For the one-dimensional problem, the instability of Flux NO is such that only results up to $t=0.2$ are shown in Table 1. From Table 1, it can be seen that the norms related to velocity are the largest and increase the fastest, which can be attributed to the poorer approximation of the $u_{y}$ and $u_{z}$ components, while $u_{x}$ is well approximated compared to these, as indicated by Figure 4. Consequently, although the training dataset for the one-dimensional problem is richer, consisting of 500 functions compared to 100 for the two-dimensional problem, it is expected to have lower stability and generalization power due to the reasons mentioned above. For the two-dimensional problem, the performance on test samples within the short-term training distribution appears to be satisfactory, as can be seen from Figure 5, 6, 14, 7 and Table 2, indicating that the solution for each component has been well resolved.
@@
Long term inference We conducted experiments over time spans significantly longer than the time domains of the functions used for training, with the initial conditions configured from the same distribution as training dataset. Due to instability issues, long-term experiments were not conducted for the one-dimensional case; instead, they were carried out only for the two-dimensional case, with the results presented in Figure 8, 18, 15, 19 and Table 3. The training functions’ time domain ranged from $t=0.25$ to $t=0.75$, with a fixed time interval of $\Delta t=0.005$, allowing for the terminal point to be reached after 100 iterations. However, in our results, using the same time intervals, the iterations extended to much longer times: $t=1.25$ (200 iterations), $t=1.5$ (250 iterations), $t=1.75$ (300 iterations), and $t=2.0$ (350 iterations). As shown in Figures 14 and 15, the error increases almost linearly with the number of iterations, affecting both the global and local characteristics of the solution and leading to progressively increasing discrepancies with the reference, as qualitatively evident in Figures 8, 18 and 19. A notable aspect of the long-term inference for the two-dimensional problem is that, unlike the one-dimensional case, even with many more iterations, the solutions output by Flux NO did not blow up.
@@
High resolution inference We conducted tests on samples at a higher resolution (96 by 96) than the resolution of the training data (64 by 64). Initial conditions were sampled from a Gaussian random field with the same power spectrum as the training data, and references were computed using the WENO-Z scheme with each snapshot spaced by a $\Delta t=0.005$. The test results are presented in Figures 9, 20, 21 and 16; trends in errors, which showed no significant difference from the original resolution, have been omitted. As seen in Figures 9 and 20, while Flux NO captures the global characteristics of the solution, local instabilities gradually appear in the snapshots at $t=1.5$ and $t=2.0$. This instability, similar to that observed in the one-dimensional case, appears to arise from the higher resolution and is somewhat related to the CFL condition. Attempts to more easily satisfy the CFL condition by reducing the interval to $\Delta t=0.0025$ for denser inference were made, yet instabilities still formed. From these experimental results, we hypothesize that as the resolution of vectorized functions increases, the domain dimension of their data distribution expands, making it challenging to approximate the data distribution itself, leading to these outcomes.